Solve the equation. You must show every step. This includes distributing, adding and subtracting a number from both sides of the equation, multipliying a number on both sides of the equation, and dividing a number on both sides of the equation.
5x + 1 = 1
______ ___
8 2

Answer by checkley77(12844)


5x + 1 = 1
______ ___
8 2
(5X+1)/8=1/2 CROSS MULTIPLY.
2(5X+1)=8*1 MULTIPLY.
10X+2=8 SUBTRACT 2.
10X=8-2 COMBINE THE INTEGERS.
10X=6 DIVIDE BY 10.
X=6/10 REDUCE THE FRACTION.
X=3/5 OR 0.6 ANS.